# Global Public Sentiment on Decentralized Finance: A Spatiotemporal Analysis of Geo-tagged Tweets from 150 Countries

### v1

@@ VI-A Limitations
Country GDP_per_capita Tweet_proportion Sentiment_score
China 9538.005 0.003 0.673
South Korea 31038.330 0.003 0.688
Cluster 1 (mean) 3075.595 0.497 0.706
Cluster 2 (mean) 4222.720 0.102 0.651
Cluster 3 (mean) 33879.133 0.280 0.665
@@ VII Conclusion
Research Question: While extensive research has explored the relationship between blockchain and sustainable economy [62, 63], our work uniquely contributes to the intersection of geospatial analysis and the blockchain ecosystem. This approach paves the way for future studies that extend the analysis to other critical social and economic issues, such as sustainability (ESG factors), unemployment, inflation, and poverty reduction, by integrating additional relevant data sources, including remote sensing and World Bank data. By emphasizing the social implications and functionalities of blockchain, our research underscores its potential to address and resolve existing social challenges and suggests avenues for improving decentralized finance (DeFi) to enhance sustainability.

### v2

@@ VI-A Limitations
Country/Cluster GDP per Capita (USD) Tweet Proportion (%) Sentiment Score
China 9,538 0.003 0.673
South Korea 31,038 0.003 0.688
Cluster 1 (mean) 3,075 0.497 0.706
Cluster 2 (mean) 4,222 0.102 0.651
Cluster 3 (mean) 33,879 0.280 0.665
Note: This table compares the GDP per capita, tweet proportion, and sentiment score of China and South Korea with the mean values of the three identified clusters. China and South Korea were identified as outliers in our analysis, as visualized in Figure 7 and further compared using box plots in Figure IV. China’s geotagged tweet data may not accurately reflect public sentiment and attention due to the country’s Twitter ban. While some individuals access Twitter through VPNs (Virtual Private Networks), the vast majority of Chinese citizens have limited knowledge of or access to VPN services. Consequently, the sentiment and public attention data for China likely represent a small, tech-savvy, and potentially non-representative subset of the population, which may introduce bias in interpreting decentralized finance discussions. South Korea, on the other hand, has unrestricted Twitter access, making its tweet data more reflective of general public opinion. However, despite its high GDP per capita, its tweet proportion remains significantly lower than those in the highest-income cluster (Cluster 3). This suggests that factors beyond economic development, such as cultural or regulatory influences, may shape decentralized finance engagement in South Korea. Overall, the comparison highlights the complexities in using geotagged tweets as a proxy for decentralized finance discussions, particularly in countries with restricted social media environments.
@@ VII Conclusion
Research Question: While extensive research has examined the relationship between blockchain and a sustainable economy [71, 72], this study uniquely contributes to the intersection of geospatial analysis and the blockchain ecosystem. Given that geospatial locations are strongly correlated with cultural [73], economic, and regulatory contexts, future research should explore how cultural factors influence public engagement with blockchain technologies and decentralized finance (DeFi). Incorporating cultural dimensions [74]—such as societal trust in technology, financial behaviors, and attitudes toward decentralization—could provide deeper insights into regional variations in blockchain adoption. Additionally, future studies could extend this analysis to other pressing social and economic issues, including sustainability (ESG factors), unemployment, inflation, and poverty reduction, by integrating additional data sources such as remote sensing and World Bank datasets. A more comprehensive understanding of these interconnections would enhance the development of blockchain-based solutions tailored to diverse geographic and cultural contexts, ultimately improving their accessibility, effectiveness, and long-term sustainability.